### v2

# Tuning of Josephson junctions - the effects of depinning physics

Abstract
Building more powerful quantum computers requires manufacturing processes with tight tolerances. To improve the tolerances on Josephson junctions, techniques to fine tune their properties after fabrication have been developed. Understanding how tuning techniques may physically modify the tunnel barrier of a Josephson junction is important and will enable these techniques to be optimised. We develop a model of junction tuning based on depinning theory to interpret a phase diagram of tuning rate. We extract the dependence on temperature, time-varying voltages and oscillation frequency. Using depinning theory we are able to show both why time-varying annealing potentials result in controlled junction tuning and how such protocols can be optimised. We examine how tuning changes the electrical breakdown of barriers and discrepancies between modeled and measured higher energy levels of transmon qubits.
Quantum computers with high qubit count, low error rates and error correcting protocols promise new computational capabilities. Processors comprised of superconducting qubits are a leading platform for utility scale quantum computation [1, 2, 3] but building these at scale and with high quality poses many challenges, several of which lie in the processor manufacture. One such component which poses manufacturing difficulties is the Josephson junction (JJ). These are most often based on stacks of Al/AlOx/Al and are typically made by shadow angle evaporation [4] where a bottom layer of thin-film aluminium is oxidised before being coated by a second layer of aluminium. Uncontrolled variation in as-fabricated JJs currently limit qubit frequency assignment accuracy to 1 % [5, 6, 7], insufficiently precise for high fidelity 2Q gates in some gate modalities [8, 9]. Not only are junctions challenging to make with tight tolerance, but their properties change over time after their manufacture in a process known as aging [10] which complicates processor design. In order to address this variation a family of techniques which allow post-manufacture parameter adjustment have been developed including modifying the junction by laser irradiation [9, 11], electron beam irradiation [12] and alternating bias assisted annealing (ABAA) [13].
In this article we introduce a refined bias protocol using a lock-in amplifier, which naturally incorporates alternating bias [13]. The lock-in allows for fast measurements of resistance at variable frequencies allowing precise resistance targetting even with large tuning speed. We explore the voltage-temperature-frequency space of the tuning process and interpret the resultant phase diagram with a depinning model. We find good agreement between this model and the three-dimensional phase diagram. We compare tuned and untuned junctions measuring tunneling IV curves, breakdown voltages and Josephson harmonics [14].
We build a setup shown schematically in Fig. 1 (a) where we connect a JJ in series with a load resistor, creating a potential divider, and connect this circuit between the output ports of a Zurich Instruments MFLI lock-in amplifier. Measuring the voltage drop across either of these elements allows us to infer the resistance of the JJ, the property we optimise in this process. We can control both the amplitude and the frequency of the AC voltage output from the lock-in amplifier. The sample containing JJs is placed on a Peltier heating stage to control the temperature of the JJ during this process.
We measure the resistance of the JJ as a function of the lock-in frequency (Fig. 1 b) and find that we get constant values up to a frequency of 8 kHz, when the measured resistance increases. This increase in resistance is a measurement artefact due to RC filtering based on a computed cut-off frequency of 40 kHz, with most capacitance arising from coaxial cabling 111Using a series resistance of 15 k and a typical capacitance for BNC cables of 80pF/m. Considering 3m of cables, we compute an RC frequency of 44 kHz.. We confine our frequency operating window to frequencies of 1 kHz and below to ensure we infer the correct resistances of the JJs.
In Fig. 1 (c) we show how resistance changes during the tuning protocol for three frequencies; 8 Hz, 81 Hz and 811 Hz. We fit these tuning curves phenomenologically to a model where the resistance increases logarithmically with time
where is the resistance at the start of the protocol and both and are fit parameters, with relating to the speed of tuning. We find this model fits all measured tuning curves well, and that tuning increases in speed with decreasing frequency. On some occasions, not shown in Fig. 1, the junction fails during the tuning process, becoming a short circuit. The rate of junction failures for a given process is another parameter we explore.
We show that the behaviour (speed and failure rate) of junctions undergoing voltage tuning can be understood qualitatively in terms of depinning theory. Depinning theory applies to a wide range of systems in which there is an interplay between disorder, temperature and applied stresses. This is seen in domain walls in magnets, the movement of vortices in superconductors, charge movement in Josephson junction arrays, dislocation movement, crack formation and even the formation of avalanches in sand piles [16, 17, 18, 19, 20].
The classic picture of depinning theory as applied to disordered media invokes two phases, a pinned static phase where the system is in a relaxed local minima and a sliding phase where local regions or domains within the system have a nonzero average velocity [21, 22]. In the pinned phase, any static or oscillating field can induce local microscopic changes, which can act as a dissipation channel [23], but not large scale movement or restructure. Above some threshold applied field (the ‘depinning threshold’) the movement of these local domains has some finite average velocity and can ultimately lead to avalanche behaviour.
We consider the junction dielectric to host domains which are trapped local minima. Applying sufficient field or temperature can move these domains, surmounting some effective potential barrier. Following Nattermann et al. [22], we define the depinning temperature where the thermal energy is equivalent to the barrier height. The depinning voltage is the point at which there is sufficient applied field for domains to continuously move through the material, resulting in an avalanche.
The regions of interest in a generic depinning theory are illustrated in Fig. 2 (a). The linear response or ‘running’ region (blue shaded area) applies for constant fields greater than the temperature dependent depinning voltage . Any constant applied field less than this will not result in a significant change in the structure of the material, it will be ‘pinned’ (the red shaded area). The zero temperature limit of is the threshold voltage .
The general depinning threshold () which separates these two regions can be defined for temperature and voltage by the relation
where is a constant which depends on the effective dimensionality of the elastic object (for 1-2 dimensional domains in a 3D medium, it can be approximated to 1) [22, 24].
When an oscillatory field is applied, a 3rd regime is formed in the phase diagram. The so-called ‘creep’ region (green region of Fig. 2 a) is the region in which domains can most easily move under application of an oscillatory field. In the creep region, the oscillating field results in hysteretic behaviour [25, 17].
A general expression for the creep threshold is difficult to derive without an explicit microscopic model. Nattermann et al. suggested [22, 24] this boundary can be found by modifying Eq. 2, replacing in the adiabatic limit. In this case depends on the oscillation frequency and the characteristic timescale required for a domain to overcome the barrier to move from its local equilibrium. In what follows, we find this significantly overestimates the shift due to variation in for sensible estimates of , suggesting a more rigorous treatment is required that goes beyond the adiabatic approximation.
Interpreting the junction tuning results in terms of depinning suggests a) if the creep region is larger for lower frequencies, and b) in the creep regime, there is a linear movement of defects without leading to full avalanche. This explains why lower frequencies result in greater resistance shifts for fixed values of temperature and voltage amplitude. At some low-frequency-limit we anticipate this trend stops, as domain movement saturates before the field polarity is reversed presenting a new energy landscape to the mobile domains. We can also understand the junction breakdown in terms of the depinning avalanche process for , which is both temperature and local microstate dependent.
To explore our interpretation we tune junctions with varying voltage and temperature at frequencies of both 1 kHz and 103 Hz. For each parameter set we tune at least three junctions. We fit each of these tuning curves to Eq. 1 and extract the speed parameter . We also compute how many junctions fail prior to 150 s of tuning process. We present this data as a function of temperature and voltage amplitude, Fig. 2 (b). A clear boundary is observed at a tuning voltage of V. This boundary is matched by DC measurements of breakdown voltage performed at varying temperature. We therefore associate this boundary with the depinning threshold and fit the DC breakdown to Eq. 2.
We continue to map out the phase diagram by measuring the tuning ‘speed’ parameter, in Eq. 1. Operating at 1 kHz, we vary the temperature and voltage amplitude of the tuning process. In Fig. 2 (c) we plot , finding a similar boundary, now at a lower voltage, consistent with the expected behaviour of . This process is then repeated at a different tuning frequency (103 Hz) with data shown in Fig. 2 (d). Again showing the expected increase in the creep regime with reducing frequency.
To fit the boundaries between regimes quantitatively is more difficult, as it requires estimates for , , and and a better understanding of the relationship between depinning velocity and the tuning speed parameter. We can estimate from junction breakdown measurements at a given temperature. Some indications for the depinning temperature can be given by observations of where the rate of thermal aging starts to increase. To give an indication of likely temperature scales, in [26, 27, 28] changes are first seen C and large scale junction modification is observed at C in [29, 30, 31, 32].
Using the approach [22, 24] given in Eq. 2, along with a finite value of , we compare to the linear boundaries shown in Fig. 2c) and d). We find that while the direction of the experimental shift with bias frequency is expected, its magnitude is actually much smaller than depinning theory would suggest (or corresponds to unrealistically small – see supplementaary material). This suggests that the dynamics in the creep regime are more complicated than in previous studies of depinning.
The phase diagram in Fig. 2 (a) indicates sharp lines demarking regions of the phase diagram. However, increases smoothly with increasing voltage, starting at 0. Likewise, the onset of junction failure starts with a small subset of junctions failing, increasing to all junctions failing at higher voltage. Operationally these boundaries are ‘fuzzy’ which is consistent with the depinning process being dependent on the specific microstate of the amorphous oxide. As a result, considerations weighing process time against yield can be used to select a working point - even when the details of the underlying mechanism are not well understood.
Mapping out this phase diagram we have identified room temperature working points with appreciable process speed as in Ref. [33]. This offers multiple advantages, key among them is that junctions which are not tuned undergo no heating and associated aging, a significant benefit for processor design. The exact mechanism of the tuning, as well as the nature of the domains being depinned, is currently unknown. More detailed mapping of the phase diagram, combined with the statistics of many junctions, will provide invaluable data to compare to potential models of junction tuning.
While the depinning picture provides a qualitative description of the observed dynamics, a quantitative model is needed to understanding the underlying microscopic processes. Time-dependent dielectric breakdown has been extensively studied in silica-based dielectrics [34, 35, 36], in magnetic tunnel junctions [37] and low-k dielectrics [38, 39], as well as AlOx films [40, 41, 42, 43]. In these materials, a variety of breakdown mechanisms have been studied, including charge trapping within the barrier, time modulation of this charge and electric field-induced distortion of bonds in the barrier. In particular studies of copper ion migration [38, 39] have considered the influence of both convective and diffusive flux of ions. These correspond to diffusion driven by the electric field itself, or by the stoichiometry gradient respectively. This suggests a possible two-process mechanism for breakdown and tuning of Josephson junctions. Namely that breakdown is enabled by the convective movement of atoms in the direction of the applied field, while the tuning of the resistance at lower fields is related to the redistribution of oxygen or defects in the barrier via lateral diffusion - mediated by the oscillating field. To test this possibility, we investigate the average barrier characteristics before and after tuning.
Following our recent work [44] we measure the electrical breakdown of junctions by ramping a DC voltage across these junctions from 0 to 1.7 V recording the current that flows through each junction as a function of voltage (Fig. 3 a). We fit the linear portion of this IV to extract the resistance of the JJ and fit the full IV to the Simmons model [45] fixing the area to that found by AFM measurements 8.86 nm2. At some voltage the junction fails and becomes short circuit which is recorded as the breakdown voltage (black star in Fig. 3 a). We compare two sets of junctions, distinguished by their tuning, with resistances shown in Fig. 3 (b). The first set of 20 junctions is tuned at 85 ∘C with a voltage of 0.92 V targeting 30 % total tuning and find an average tuning of 34.6 0.8 %. This systematic overshoot occurs due to continued resistance changes over time, after we have hit our target value, as shown in Ref. 33. During this process two junctions failed. The second set of 25 junctions are fabricated on the same chip, undergo the same 85 ∘C heating as the tuned junctions, but are not tuned by an alternating voltage. Despite undergoing no electrical tuning they are 4 % higher in resistance than the junctions before tuning due to the thermal processing they undergo, indicative of why a room temperature process is advantageous.
In Fig. 3 (c) we see a small increase to the average fitted barrier thickness with distributions mostly overlapping for tuned and untuned junctions. In Fig. 3 (d) we see a clear increase in the barrier height for the junctions which have been tuned relative to those that have not. These results indicate that the tuning process modifies the atomic structure of the barrier more substantially than its thickness.
In Fig. 3 (e) we show that on average junctions which have undergone tuning have a larger breakdown voltage, increasing by 5 %. However, despite the increase in average breakdown voltage, the lowest breakdown voltage recorded was from a junction that had undergone tuning. The breakdown voltage is given by
where is the local thickness of the barrier and is the local dielectric strength of the barrier. Both of these values may, in principle, vary across the barrier [46, 47] and be modified by tuning. In the depinning model we describe above domains are mobile under the application of electric fields. These fields will be the strongest at the thinnest regions of the barrier which means that the the creep rate will also be greatest at these thinnest points. Such weak points in the barrier have been suggested to contribute to non-idealities [47, 14, 48] and breakdown [42] of the JJs. An approach that reduces the non-idealities associated to pin holes or weak points may have advantages in homogenising the current phase relations of junctions.
The higher-energy states of transmon qubits have recently been suggested to be sensitive to the nanostructure of the barrier in JJs [14]. In Fig. 3 (f) we show the results of measuring two coaxmon qubits [49] where one qubit has had a junction tuned by 34 % and the other has experienced the same thermal treatment but not been tuned. The results of Ref. 14 would suggest that by substantially modifying the barrier we might expect to see some difference to the frequencies of these higher levels. We follow the protocol from Ref. 50, treating the coaxmon as a qudit and tuning up a series of gates across the 6 energy levels shown schematically in Fig. 3 (f). We compute the Josephson energy and charging energy using qutip [51] by considering the first two transition frequencies and constraining them to match the results from the transmon Hamiltonian
where and are the charge and phase operators respectively and and are the charging and Josephson energies [52]. The tuned (untuned) coaxmon have junction resistances of 5802 (6021 ), charging energies of 166 MHz (168 MHz) and Josephson energies of 23.2 GHz (22.2 GHz). In Fig. 3 (f) we show the difference between the measured transition frequencies and the computed transition frequencies for the different qudit states for both a tuned and untuned coaxmon. We find there is no substantial difference between the two coaxmons.
By mapping out the voltage-temperature-frequency phase diagram of tuning under oscillatory bias we have shown that the tuning process is qualitatively described by depinning theory. The domains within the amorphous oxide barrier creep between local minima driven by both temperature and the oscillating field. Depinning theory by itself does not point to a particular origin of these local domains, as it is consistent with small scale defects [53], movement of grain boundaries [54] or other microscopic models [23]. Nonetheless by interpreting the results in terms of depinning theory, the working point can be chosen to improve the junction tuning protocol, choosing how deeply into the creep-region one wishes to operate. We find a set of parameters that allow junction tuning at room temperature with appreciable speed, where untuned junctions are also not subject to heating and associated modifications.
We show that tuning modifies, not only the resistance of the junctions, but also the barrier height in the junction and the breakdown voltage of the junctions. We contrast this to a measurement of ‘Josephson Harmonics’ which appear unchanged comparing tuned and untuned junctions. The interpretation in Ref. 14 suggests that deviations in these harmonics can be ascribed to the thinnest points of junctions. Taken together our measurements suggest that the depinning process modifying the junction resistance is not simply ‘filling in’ the thinnest points of barriers, but fundamentally modifies the junction microstructure.
Acknowledgements.
We extend our thanks to the OQC Fabrication team for discussions relating to this work, the fabrication of junctions and AFM measurements of junction area. We acknowledge the OQC Hardware team for developments across the stack which enabled these measurements. The authors thank Dave Pappas and Simon Brown for useful discussions. The authors thank Peter Leek and Jonathan Burnett for reviewing this manuscript. Computational resources were provided by the Australian National Computational Infrastructure facility (NCI). We thank the Royal Holloway University of London SuperFab Facility for their support.References
- Arute et al. [2019] F. Arute, K. Arya, R. Babbush, D. Bacon, J. C. Bardin, R. Barends, R. Biswas, S. Boixo, F. G. Brandao, D. A. Buell, et al., Quantum supremacy using a programmable superconducting processor, Nature 574, 505 (2019).
- Chow et al. [2021] J. Chow, O. Dial, and J. Gambetta, Ibm quantum breaks the 100-qubit processor barrier, IBM Research Blog 2 (2021).
- AI [2023] G. Q. AI, Suppressing quantum errors by scaling a surface code logical qubit, Nature 614, 676 (2023).
- Kreikebaum et al. [2020] J. Kreikebaum, K. O’Brien, A. Morvan, and I. Siddiqi, Improving wafer-scale josephson junction resistance variation in superconducting quantum coherent circuits, Superconductor Science and Technology 33, 06LT02 (2020).
- Osman et al. [2023] A. Osman, J. Fernández-Pendás, C. Warren, S. Kosen, M. Scigliuzzo, A. Frisk Kockum, G. Tancredi, A. Fadavi Roudsari, and J. Bylander, Mitigation of frequency collisions in superconducting quantum processors, Physical Review Research 5, 043001 (2023).
- Acharya et al. [2024] N. Acharya, R. Armstrong, Y. Balaji, K. G. Crawford, J. C. Gates, P. C. Gow, O. W. Kennedy, R. D. Pothuraju, K. Shahbazi, and C. D. Shelly, Integration of through-sapphire substrate machining with superconducting quantum processors, Advanced Materials , 2411780 (2024).
- Van Damme et al. [2024] J. Van Damme, S. Massar, R. Acharya, T. Ivanov, D. Perez Lozano, Y. Canvel, M. Demarets, D. Vangoidsenhoven, Y. Hermans, J. Lai, et al., Advanced cmos manufacturing of superconducting qubits on 300 mm wafers, Nature , 1 (2024).
- Morvan et al. [2022] A. Morvan, L. Chen, J. M. Larson, D. I. Santiago, and I. Siddiqi, Optimizing frequency allocation for fixed-frequency superconducting quantum processors, Physical Review Research 4, 023079 (2022).
- Hertzberg et al. [2021] J. B. Hertzberg, E. J. Zhang, S. Rosenblatt, E. Magesan, J. A. Smolin, J.-B. Yau, V. P. Adiga, M. Sandberg, M. Brink, J. M. Chow, et al., Laser-annealing josephson junctions for yielding scaled-up superconducting quantum processors, npj Quantum Information 7, 129 (2021).
- Koppinen et al. [2007a] P. Koppinen, L. Väistö, and I. Maasilta, Complete stabilization and improvement of the characteristics of tunnel junctions by thermal annealing, Applied physics letters 90 (2007a).
- Kim et al. [2022] H. Kim, C. Jünger, A. Morvan, E. S. Barnard, W. P. Livingston, M. Altoé, Y. Kim, C. Song, L. Chen, J. M. Kreikebaum, et al., Effects of laser-annealing on fixed-frequency superconducting qubits, Applied Physics Letters 121 (2022).
- Balaji et al. [2024] Y. Balaji, N. Acharya, R. Armstrong, K. G. Crawford, S. Danilin, T. Dixon, O. W. Kennedy, R. D. Pothuraju, K. Shahbazi, and C. D. Shelly, Electron-beam annealing of josephson junctions for frequency tuning of quantum processors, arXiv preprint arXiv:2402.17395 (2024).
- Pappas et al. [2024] D. P. Pappas, M. Field, C. J. Kopas, J. A. Howard, X. Wang, E. Lachman, J. Oh, L. Zhou, A. Gold, G. M. Stiehl, et al., Alternating-bias assisted annealing of amorphous oxide tunnel junctions, Communications Materials 5, 150 (2024).
- Willsch et al. [2024] D. Willsch, D. Rieger, P. Winkel, M. Willsch, C. Dickel, J. Krause, Y. Ando, R. Lescanne, Z. Leghtas, N. T. Bronn, et al., Observation of josephson harmonics in tunnel junctions, Nature Physics , 1 (2024).
- Note [1] Using a series resistance of 15 k and a typical capacitance for BNC cables of 80pF/m. Considering 3m of cables, we compute an RC frequency of 44 kHz.
- Wiese [2022] K. J. Wiese, Theory and experiments for disordered elastic manifolds, depinning, avalanches, and sandpiles, Reports on Progress in Physics 85, 086502 (2022).
- Brazovskii and Nattermann [2004] S. Brazovskii and T. Nattermann, Pinning and sliding of driven elastic systems: from domain walls to charge density waves, Advances in Physics 53, 177 (2004).
- Vogt et al. [2014] N. Vogt, R. Schäfer, H. Rotzinger, W. Cui, A. Fiebig, A. Shnirman, and A. V. Ustinov, One-dimensional Josephson junction arrays: Lifting the Coulomb blockade by depinning, Physical Review B 92 (2014).
- Vogt et al. [2016] N. Vogt, J. H. Cole, and A. Shnirman, De-pinning of disordered bosonic chains, New Journal of Physics 18, 053026 (2016).
- Cedergren et al. [2017] K. Cedergren, R. Ackroyd, S. Kafanov, N. Vogt, A. Shnirman, and T. Duty, Insulating Josephson Junction Chains as Pinned Luttinger Liquids, Physical Review Letters 119, 167701 (2017).
- Middleton and Fisher [1991] A. A. Middleton and D. S. Fisher, Critical behavior of pinned charge-density waves below the threshold for sliding, Physical Review Letters 66, 92 (1991).
- Nattermann et al. [2001] T. Nattermann, V. Pokrovsky, and V. M. Vinokur, Hysteretic Dynamics of Domain Walls at Finite Temperatures, Physical Review Letters 87, 197005 (2001).
- Müller et al. [2019] C. Müller, J. H. Cole, and J. Lisenfeld, Towards understanding two-level-systems in amorphous solids: insights from quantum circuits, Reports on Progress in Physics 82, 124501 (2019).
- Nattermann and Pokrovsky [2004] T. Nattermann and V. Pokrovsky, Hysteresis mediated by a domain wall motion, Physica A 340, 625 (2004).
- Schütze and Nattermann [2011] F. Schütze and T. Nattermann, Mean-field theory for driven domain walls in disordered environments, Physical Review B 83, 024412 (2011).
- Imamura et al. [1992] T. Imamura, T. Shiota, and S. Hasuo, Fabrication of high quality nb/alo/sub x/-al/nb josephson junctions. i. sputtered nb films for junction electrodes, IEEE Transactions on Applied Superconductivity 2, 1 (1992).
- Oliva and Monaco [1994] A. Oliva and R. Monaco, Annealing properties of high quality nb/al-alo/sub x//nb tunnel junctions, IEEE Transactions on Applied Superconductivity 4, 25 (1994).
- Migacz and Huber [2003] J. Migacz and M. Huber, Thermal annealing of nb/al-alo/sub x//nb josephson junctions, IEEE Transactions on Applied Superconductivity 13, 123 (2003).
- Pop et al. [2012] I. M. Pop, T. Fournier, T. Crozes, F. Lecocq, I. Matei, B. Pannetier, O. Buisson, and W. Guichard, Fabrication of stable and reproducible submicron tunnel junctions, Journal of Vacuum Science & Technology B 30 (2012).
- Julin et al. [2010] J. K. Julin, P. J. Koppinen, and I. J. Maasilta, Reduction of low-frequency 1/f noise in al–alox–al tunnel junctions by thermal annealing, Applied Physics Letters 97, 152501 (2010).
- Koppinen et al. [2007b] P. J. Koppinen, L. M. Väistö, and I. J. Maasilta, Complete stabilization and improvement of the characteristics of tunnel junctions by thermal annealing, Applied Physics Letters 90, 053503 (2007b).
- Korshakov et al. [2024] N. D. Korshakov, D. O. Moskalev, A. A. Soloviova, D. A. Moskaleva, E. S. Lotkov, A. R. Ibragimov, M. V. Androschuk, I. A. Ryzhikov, Y. V. Panfilov, and I. A. Rodionov, Aluminum josephson junction microstructure and electrical properties modification with thermal annealing, arXiv preprint arXiv:2403.02179 (2024).
- Wang et al. [2024a] X. Wang, J. Howard, E. A. Sete, G. Stiehl, C. Kopas, S. Poletto, X. Wu, M. Field, N. Sharac, C. Eckberg, et al., Precision frequency tuning of tunable transmon qubits using alternating-bias assisted annealing, in 2024 IEEE International Conference on Quantum Computing and Engineering (QCE), Vol. 1 (IEEE, 2024) pp. 1315–1323.
- McPherson [2012] J. W. McPherson, Time dependent dielectric breakdown physics–models revisited, Microelectronics Reliability 52, 1753 (2012).
- Moxim et al. [2022] S. J. Moxim, F. V. Sharov, D. R. Hughart, G. S. Haase, C. G. McKay, and P. M. Lenahan, Atomic-scale defects generated in the early/intermediate stages of dielectric breakdown in si/sio2 transistors, Applied Physics Letters 120 (2022).
- McPherson et al. [2003] J. McPherson, J. Kim, A. Shanware, and H. Mogul, Thermochemical description of dielectric breakdown in high dielectric constant materials, Applied Physics Letters 82, 2121 (2003).
- Amara-Dababi et al. [2012] S. Amara-Dababi, H. Bea, R. Sousa, K. Mackay, and B. Dieny, Modelling of time-dependent dielectric barrier breakdown mechanisms in mgo-based magnetic tunnel junctions, Journal of physics D: applied physics 45, 295002 (2012).
- Borja et al. [2016] J. P. Borja, T.-M. Lu, and J. Plawsky, Dielectric breakdown in gigascale electronics: Time dependent failure mechanisms (Springer, 2016).
- Borja et al. [2012] J. Borja, J. L. Plawsky, T.-M. Lu, and W. N. Gill, Impact of frequency from a bipolar applied field on dielectric breakdown for low- materials, IEEE transactions on electron devices 59, 1745 (2012).
- Schaefer et al. [2011] D. Schaefer, P. Fichtner, M. Carara, L. Schelp, and L. Dorneles, Dielectric breakdown in alox tunnelling barriers, Journal of Physics D: Applied Physics 44, 135403 (2011).
- Kolodzey et al. [2000] J. Kolodzey, E. A. Chowdhury, T. N. Adam, G. Qui, I. Rau, J. O. Olowolafe, J. S. Suehle, and Y. Chen, Electrical conduction and dielectric breakdown in aluminum oxide insulators on silicon, IEEE Transactions on Electron Devices 47, 121 (2000).
- Tolpygo and Amparo [2008] S. K. Tolpygo and D. Amparo, Electrical stress effect on Josephson tunneling through ultrathin AlOx barrier in Nb/Al/AlOx/Nb junctions, Journal of Applied Physics 104, 063904 (2008).
- De Wit et al. [1976] H. De Wit, C. Wijenberg, and C. Crevecoeur, The dielectric breakdown of anodic aluminum oxide, Journal of The Electrochemical Society 123, 1479 (1976).
- Kennedy et al. [2025] O. W. Kennedy, K. G. Crawford, K. Shahbazi, and C. D. Shelly, Analysis of josephson junction barrier variation – a combined stem, breakdown and monte-carlo approach, arXiv preprint arXiv:2502.13072 (2025).
- Simmons [1963] J. G. Simmons, Generalized formula for the electric tunnel effect between similar electrodes separated by a thin insulating film, Journal of applied physics 34, 1793 (1963).
- Cyster et al. [2020] M. J. Cyster, J. S. Smith, J. A. Vaitkus, N. Vogt, S. P. Russo, and J. H. Cole, Effect of atomic structure on the electrical response of aluminum oxide tunnel junctions, Physical Review Research 2, 013110 (2020).
- Bayros et al. [2024] K. Bayros, M. Cyster, J. Smith, and J. Cole, Influence of pinholes and weak-points in aluminum-oxide josephson junctions, Physical Review Materials 8, 046202 (2024).
- Greibe et al. [2011] T. Greibe, M. P. V. Stenberg, C. M. Wilson, T. Bauch, V. S. Shumeiko, and P. Delsing, Are ”pinholes” the cause of excess current in superconducting tunnel junctions? A study of Andreev current in highly resistive junctions, Physical Review Letters 106, 1657 (2011).
- Rahamim et al. [2017] J. Rahamim, T. Behrle, M. Peterer, A. Patterson, P. Spring, T. Tsunoda, R. Manenti, G. Tancredi, and P. Leek, Double-sided coaxial circuit qed with out-of-plane wiring, Applied Physics Letters 110 (2017).
- Wang et al. [2024b] Z. Wang, R. W. Parker, E. Champion, and M. S. Blok, Systematic study of high transmon qudits up to , arXiv preprint arXiv:2407.17407 (2024b).
- Johansson et al. [2012] J. R. Johansson, P. D. Nation, and F. Nori, Qutip: An open-source python framework for the dynamics of open quantum systems, Computer physics communications 183, 1760 (2012).
- Koch et al. [2007] J. Koch, T. M. Yu, J. Gambetta, A. A. Houck, D. I. Schuster, J. Majer, A. Blais, M. H. Devoret, S. M. Girvin, and R. J. Schoelkopf, Charge-insensitive qubit design derived from the cooper pair box, Physical Review A—Atomic, Molecular, and Optical Physics 76, 042319 (2007).
- Strand and Shluger [2024] J. Strand and A. L. Shluger, On the structure of oxygen deficient amorphous oxide films, Advanced Science 11, 2306243 (2024).
- Oh et al. [2025] J.-S. Oh, C. J. Kopas, H. Cansizoglu, J. Y. Mutus, K. Yadavalli, T.-H. Kim, M. Kramer, A. H. King, and L. Zhou, Correlating aluminum layer deposition rates, josephson junction microstructure, and superconducting qubits’ performance, Acta Materialia 284, 120631 (2025).

### v3

@@
-Tuning of Josephson junctions - the effects of depinning physics
+A Depinning Model for Josephson Junction Tuning
 Abstract
-Building more powerful quantum computers requires manufacturing processes with tight tolerances.
-To improve the tolerances on Josephson junctions, techniques to fine tune their properties after fabrication have been developed.
+Superconducting quantum computers leverage superconducting qubits, systems with well-defined energy levels, the spacing of which is determined by the properties of embedded Josephson junctions.
+Indeterminacy of Josephson junction manufacturing limits large scale superconducting circuits, however junction-by-junction tuning affords the possibility of an order of magnitude improvement of determinacy - although the origin of this tuning effect is uncertain.
 Understanding how tuning techniques may physically modify the tunnel barrier of a Josephson junction is important and will enable these techniques to be optimised.
@@
 We extract the dependence on temperature, time-varying voltages and oscillation frequency.
-Using depinning theory we are able to show both why time-varying annealing potentials result in controlled junction tuning and how such protocols can be optimised.
-We examine how tuning changes the electrical breakdown of barriers and discrepancies between modeled and measured higher energy levels of transmon qubits.
+Comparing the effect of junction tuning on room-temperature IVs, electrical breakdown and frequency of higher energy levels of transmon qubits indicates that the tuning process leaves the barrier thickness unchanged but modifies its composition.
 Quantum computers with high qubit count, low error rates and error correcting protocols promise new computational capabilities.
-Processors comprised of superconducting qubits are a leading platform for utility scale quantum computation [1, 2, 3] but building these at scale and with high quality poses many challenges, several of which lie in the processor manufacture.
-One such component which poses manufacturing difficulties is the Josephson junction (JJ).
+Processors comprised of superconducting qubits are a leading platform for utility scale quantum computation [1, 2, 3] but building these at scale and with high quality poses many challenges.
+One component which poses manufacturing difficulties is the Josephson junction (JJ).
 These are most often based on stacks of Al/AlOx/Al and are typically made by shadow angle evaporation [4] where a bottom layer of thin-film aluminium is oxidised before being coated by a second layer of aluminium.
-Uncontrolled variation in as-fabricated JJs currently limit qubit frequency assignment accuracy to 1 % [5, 6, 7], insufficiently precise for high fidelity 2Q gates in some gate modalities [8, 9].
+State-of-the-art variation in as-fabricated JJs currently limits junction resistance accuracy to 2 % [5, 6, 7, 8, 9, 10].
+The problem of junction variation will likely be exacerbated in larger wafers necessitated by larger processors and by the use of long-range couplers [11].
+Simulations show that current manufacturing accuracy is insufficient to build large, high quality, quantum processors [10, 8, 5].
+Improvements to frequency assignment enabled by tuning protocols, have been shown to improve processor yield experimentally [9].
 Not only are junctions challenging to make with tight tolerance, but their properties change over time after their manufacture in a process known as aging [12] which complicates processor design.
 In order to address this variation a family of techniques which allow post-manufacture parameter adjustment have been developed including modifying the junction by laser irradiation [5, 13], electron beam irradiation [14] and alternating bias assisted annealing (ABAA) [15].
-In this article we introduce a refined bias protocol using a lock-in amplifier, which naturally incorporates alternating bias [13].
-The lock-in allows for fast measurements of resistance at variable frequencies allowing precise resistance targetting even with large tuning speed.
-We explore the voltage-temperature-frequency space of the tuning process and interpret the resultant phase diagram with a depinning model.
-We find good agreement between this model and the three-dimensional phase diagram.
+These techniques, referred to here as tuning, cause the resistance of junctions to increase only.
+In this letter, we explore the voltage-temperature-frequency space, facilitated by a simplified lock-in tuning technique adapted from [15].
+We interpret the resultant phase diagram using depinning theory which applies to a wide range of systems in which there is an interplay between disorder, temperature and applied stresses.
+This is seen in domain walls in magnets, the movement of vortices in superconductors, charge movement in Josephson junction arrays, dislocation movement, crack formation and even the formation of avalanches in sand piles [16, 17, 18, 19, 20].
+We find good agreement between our depinning model and the three-dimensional phase diagram.
 We compare tuned and untuned junctions measuring tunneling IV curves, breakdown voltages and Josephson harmonics [21].
@@
 The sample containing JJs is placed on a Peltier heating stage to control the temperature of the JJ during this process.
+JJs are fabricated in Ref [22], Dolan bridge junctions with an area of 84,000 nm2 are fabricated by shadow evaporations before and after a static oxidation step.
+An atomic force micrograph of a typical junction is shown in the supplemental materials [23] which includes Ref.
+[24].
 We measure the resistance of the JJ as a function of the lock-in frequency (Fig.
-1 b) and find that we get constant values up to a frequency of 8 kHz, when the measured resistance increases.
-This increase in resistance is a measurement artefact due to RC filtering based on a computed cut-off frequency of 40 kHz, with most capacitance arising from coaxial cabling 111Using a series resistance of 15 k and a typical capacitance for BNC cables of 80pF/m.
-Considering 3m of cables, we compute an RC frequency of 44 kHz..
+1 b) and find that we measure a constant value up to a frequency of 8 kHz, when the measured resistance increases.
+This increase in resistance is a measurement artefact due to RC filtering based on a computed cut-off frequency of 40 kHz, with most capacitance arising from coaxial cabling - a series resistance of 15 k (10 k load and 5 k junction resistance) together with 3 m of BNC cables (typical capacitance of 80 pF/m) gives an RC frequency of 44 kHz.
 We confine our frequency operating window to frequencies of 1 kHz and below to ensure we infer the correct resistances of the JJs.
@@
 We show that the behaviour (speed and failure rate) of junctions undergoing voltage tuning can be understood qualitatively in terms of depinning theory.
-Depinning theory applies to a wide range of systems in which there is an interplay between disorder, temperature and applied stresses.
-This is seen in domain walls in magnets, the movement of vortices in superconductors, charge movement in Josephson junction arrays, dislocation movement, crack formation and even the formation of avalanches in sand piles [16, 17, 18, 19, 20].
-The classic picture of depinning theory as applied to disordered media invokes two phases, a pinned static phase where the system is in a relaxed local minima and a sliding phase where local regions or domains within the system have a nonzero average velocity [21, 22].
+The classic picture of depinning theory as applied to disordered media invokes two phases, a pinned static phase where the system is in a relaxed local minimum and a sliding phase where local regions or domains within the system have a nonzero average velocity [25, 26].
+The domains of depinning theory are a general concept, which may sometimes be instantiated by magnetic domains specifically, but this commonality in terminology is not fundamental.
+For example, domains may similarly be regions of sand grains, or any other type of effective degree of freedom that ‘depins’ in a disordered system.
 In the pinned phase, any static or oscillating field can induce local microscopic changes, which can act as a dissipation channel [27], but not large scale movement or restructure.
@@
 We consider the junction dielectric to host domains which are trapped local minima.
+These domains may be 0D vacancies/interstitials, 1D chains of such vacancies, or 2D boundary regions of lower density amorphous material or something else - the depinning treatment here is agnostic to the underlying microphysics.
 Applying sufficient field or temperature can move these domains, surmounting some effective potential barrier.
@@
 This explains why lower frequencies result in greater resistance shifts for fixed values of temperature and voltage amplitude.
-At some low-frequency-limit we anticipate this trend stops, as domain movement saturates before the field polarity is reversed presenting a new energy landscape to the mobile domains.
+At some low-frequency-limit we anticipate this trend stops, as domain movement saturates before the field polarity is reversed, presenting a new energy landscape to the mobile domains.
 We can also understand the junction breakdown in terms of the depinning avalanche process for , which is both temperature and local microstate dependent.
@@
 This process is then repeated at a different tuning frequency (103 Hz) with data shown in Fig.
-2 (d).
-Again showing the expected increase in the creep regime with reducing frequency.
+2 (d), showing the expected increase in the creep regime with reducing frequency.
 To fit the boundaries between regimes quantitatively is more difficult, as it requires estimates for , , and and a better understanding of the relationship between depinning velocity and the tuning speed parameter.
@@
 Using the approach [26, 28] given in Eq.
-2, along with a finite value of , we compare to the linear boundaries shown in Fig.
+2, along with a finite value of , we compare to the stable/creep boundaries shown in Fig.
 2c) and d).
-We find that while the direction of the experimental shift with bias frequency is expected, its magnitude is actually much smaller than depinning theory would suggest (or corresponds to unrealistically small – see supplementaary material).
+We find that while the direction of the experimental shift with bias frequency is expected, its magnitude is actually much smaller than depinning theory would suggest (or corresponds to unrealistically small – see supplementary material [23]).
 This suggests that the dynamics in the creep regime are more complicated than in previous studies of depinning.
-The phase diagram in Fig.
-2 (a) indicates sharp lines demarking regions of the phase diagram.
-However, increases smoothly with increasing voltage, starting at 0.
-Likewise, the onset of junction failure starts with a small subset of junctions failing, increasing to all junctions failing at higher voltage.
-Operationally these boundaries are ‘fuzzy’ which is consistent with the depinning process being dependent on the specific microstate of the amorphous oxide.
-As a result, considerations weighing process time against yield can be used to select a working point - even when the details of the underlying mechanism are not well understood.
+Fig.
+2 (a) indicates sharp lines demarking regions of the phase diagram, however, at this stage it is not clear that these boundaries are true phase transitions.
+The speed parameter increases smoothly with increasing tuning voltage (the data is shown with alternate visualisation in the supplementary materials [23]).
+This may be indicative of a second order phase transition between the stable and creep regimes.
 Mapping out this phase diagram we have identified room temperature working points with appreciable process speed as in Ref.
@@
 This suggests a possible two-process mechanism for breakdown and tuning of Josephson junctions.
-Namely that breakdown is enabled by the convective movement of atoms in the direction of the applied field, while the tuning of the resistance at lower fields is related to the redistribution of oxygen or defects in the barrier via lateral diffusion - mediated by the oscillating field.
+Namely that breakdown is enabled by the convective movement of atoms from within the barrier in the direction of the applied field, while the tuning of the resistance at lower fields is related to the redistribution of oxygen or defects from within the barrier via diffusion - mediated by the oscillating field.
 To test this possibility, we investigate the average barrier characteristics before and after tuning.
@@
 This systematic overshoot occurs due to continued resistance changes over time, after we have hit our target value, as shown in Ref.
-33.
+6.
 During this process two junctions failed.
 The second set of 25 junctions are fabricated on the same chip, undergo the same 85 ∘C heating as the tuned junctions, but are not tuned by an alternating voltage.
-Despite undergoing no electrical tuning they are 4 % higher in resistance than the junctions before tuning due to the thermal processing they undergo, indicative of why a room temperature process is advantageous.
+Despite undergoing no electrical tuning they are 4 % higher in resistance than the junctions before tuning due to the thermal processing they undergo.
+This shows why a room temperature process (as shown in the supplementary materials [23]) is advantageous as heating chips above storage temperatures causes uncontrolled changes in junction properties alongside the intentional tuning process.
 In Fig.
@@
 Such weak points in the barrier have been suggested to contribute to non-idealities [50, 21, 51] and breakdown [45] of the JJs.
-An approach that reduces the non-idealities associated to pin holes or weak points may have advantages in homogenising the current phase relations of junctions.
+An approach that reduces the non-idealities associated to pin holes or weak points may have advantages in homogenising the current-phase relations of junctions.
 The higher-energy states of transmon qubits have recently been suggested to be sensitive to the nanostructure of the barrier in JJs [21].
@@
 We follow the protocol from Ref.
-50, treating the coaxmon as a qudit and tuning up a series of gates across the 6 energy levels shown schematically in Fig.
+53, treating the coaxmon as a qudit and tuning up a series of gates across the 6 energy levels shown schematically in Fig.
 3 (f).
@@
 3 (f) we show the difference between the measured transition frequencies and the computed transition frequencies for the different qudit states for both a tuned and untuned coaxmon.
-We find there is no substantial difference between the two coaxmons.
+We find the relative difference in model/experiment discrepancies for tuned and untuned junctions to be small relative to lab-to-lab variation or even qubit-to-qubit variation from a single institute presented in Ref.
+[21] and therefore conclude that the tuning process has not substantially modified the Josephson harmonics in this instance.
+Statistical studies comparing multiple tuned and untuned qubits may allow small effects to be resolved in future studies.
 By mapping out the voltage-temperature-frequency phase diagram of tuning under oscillatory bias we have shown that the tuning process is qualitatively described by depinning theory.
@@
 Siddiqi, Improving wafer-scale josephson junction resistance variation in superconducting quantum coherent circuits, Superconductor Science and Technology 33, 06LT02 (2020).
-- Osman et al.
-[2023] A.
-Osman, J.
-Fernández-Pendás, C.
-Warren, S.
-Kosen, M.
-Scigliuzzo, A.
-Frisk Kockum, G.
-Tancredi, A.
-Fadavi Roudsari, and J.
-Bylander, Mitigation of frequency collisions in superconducting quantum processors, Physical Review Research 5, 043001 (2023).
-- Acharya et al.
-[2024] N.
-Acharya, R.
-Armstrong, Y.
-Balaji, K.
-G.
-Crawford, J.
-C.
-Gates, P.
-C.
-Gow, O.
-W.
-Kennedy, R.
-D.
-Pothuraju, K.
-Shahbazi, and C.
-D.
-Shelly, Integration of through-sapphire substrate machining with superconducting quantum processors, Advanced Materials , 2411780 (2024).
-- Van Damme et al.
-[2024] J.
-Van Damme, S.
-Massar, R.
-Acharya, T.
-Ivanov, D.
-Perez Lozano, Y.
-Canvel, M.
-Demarets, D.
-Vangoidsenhoven, Y.
-Hermans, J.
-Lai, et al., Advanced cmos manufacturing of superconducting qubits on 300 mm wafers, Nature , 1 (2024).
-- Morvan et al.
-[2022] A.
-Morvan, L.
-Chen, J.
-M.
-Larson, D.
-I.
-Santiago, and I.
-Siddiqi, Optimizing frequency allocation for fixed-frequency superconducting quantum processors, Physical Review Research 4, 023079 (2022).
 - Hertzberg et al.
@@
 Chow, et al., Laser-annealing josephson junctions for yielding scaled-up superconducting quantum processors, npj Quantum Information 7, 129 (2021).
+- Wang et al.
+[2024a] X.
+Wang, J.
+Howard, E.
+A.
+Sete, G.
+Stiehl, C.
+Kopas, S.
+Poletto, X.
+Wu, M.
+Field, N.
+Sharac, C.
+Eckberg, et al., Precision frequency tuning of tunable transmon qubits using alternating-bias assisted annealing, in 2024 IEEE International Conference on Quantum Computing and Engineering (QCE), Vol.
+1 (IEEE, 2024) pp.
+1315–1323.
+- Muthusubramanian et al.
+[2024] N.
+Muthusubramanian, M.
+Finkel, P.
+Duivestein, C.
+Zachariadis, S.
+L.
+van der Meer, H.
+M.
+Veen, M.
+W.
+Beekman, T.
+Stavenga, A.
+Bruno, and L.
+DiCarlo, Wafer-scale uniformity of dolan-bridge and bridgeless manhattan-style josephson junctions for superconducting quantum processors, Quantum Science and Technology 9, 025006 (2024).
+- Osman et al.
+[2023] A.
+Osman, J.
+Fernández-Pendás, C.
+Warren, S.
+Kosen, M.
+Scigliuzzo, A.
+Frisk Kockum, G.
+Tancredi, A.
+Fadavi Roudsari, and J.
+Bylander, Mitigation of frequency collisions in superconducting quantum processors, Physical Review Research 5, 043001 (2023).
+- Zhang et al.
+[2022] E.
+J.
+Zhang, S.
+Srinivasan, N.
+Sundaresan, D.
+F.
+Bogorin, Y.
+Martin, J.
+B.
+Hertzberg, J.
+Timmerwilke, E.
+J.
+Pritchett, J.-B.
+Yau, C.
+Wang, et al., High-performance superconducting quantum processors via laser annealing of transmon qubits, Science Advances 8, eabi6690 (2022).
+- Morvan et al.
+[2022] A.
+Morvan, L.
+Chen, J.
+M.
+Larson, D.
+I.
+Santiago, and I.
+Siddiqi, Optimizing frequency allocation for fixed-frequency superconducting quantum processors, Physical Review Research 4, 023079 (2022).
+- Du et al. [2025] Z. Du, S. Kan, S. Stein, Z. Liang, A. Li, and Y. Mao, Hardware-aware compilation for chip-to-chip coupler-connected modular quantum systems, arXiv preprint arXiv:2505.09036 (2025).
 - Koppinen et al.
@@
 Stiehl, et al., Alternating-bias assisted annealing of amorphous oxide tunnel junctions, Communications Materials 5, 150 (2024).
-- Willsch et al.
-[2024] D.
-Willsch, D.
-Rieger, P.
-Winkel, M.
-Willsch, C.
-Dickel, J.
-Krause, Y.
-Ando, R.
-Lescanne, Z.
-Leghtas, N.
-T.
-Bronn, et al., Observation of josephson harmonics in tunnel junctions, Nature Physics , 1 (2024).
-- Note [1] Using a series resistance of 15 k and a typical capacitance for BNC cables of 80pF/m. Considering 3m of cables, we compute an RC frequency of 44 kHz.
 - Wiese [2022] K. J. Wiese, Theory and experiments for disordered elastic manifolds, depinning, avalanches, and sandpiles, Reports on Progress in Physics 85, 086502 (2022).
@@
 Duty, Insulating Josephson Junction Chains as Pinned Luttinger Liquids, Physical Review Letters 119, 167701 (2017).
+- Willsch et al.
+[2024] D.
+Willsch, D.
+Rieger, P.
+Winkel, M.
+Willsch, C.
+Dickel, J.
+Krause, Y.
+Ando, R.
+Lescanne, Z.
+Leghtas, N.
+T.
+Bronn, et al., Observation of josephson harmonics in tunnel junctions, Nature Physics , 1 (2024).
+- Acharya et al.
+[2024] N.
+Acharya, R.
+Armstrong, Y.
+Balaji, K.
+G.
+Crawford, J.
+C.
+Gates, P.
+C.
+Gow, O.
+W.
+Kennedy, R.
+D.
+Pothuraju, K.
+Shahbazi, and C.
+D.
+Shelly, Integration of through-sapphire substrate machining with superconducting quantum processors, Advanced Materials , 2411780 (2024).
+- [23] See Supplemental Material at URL-will-be-inserted-by-publisher.
+- Dobrovinskaya et al. [2009] E. R. Dobrovinskaya, L. A. Lytvynov, and V. Pishchik, Properties of sapphire, Sapphire , 55 (2009).
 - Middleton and Fisher [1991] A. A. Middleton and D. S. Fisher, Critical behavior of pinned charge-density waves below the threshold for sliding, Physical Review Letters 66, 92 (1991).
@@
 Rodionov, Aluminum josephson junction microstructure and electrical properties modification with thermal annealing, arXiv preprint arXiv:2403.02179 (2024).
-- Wang et al.
-[2024a] X.
-Wang, J.
-Howard, E.
-A.
-Sete, G.
-Stiehl, C.
-Kopas, S.
-Poletto, X.
-Wu, M.
-Field, N.
-Sharac, C.
-Eckberg, et al., Precision frequency tuning of tunable transmon qubits using alternating-bias assisted annealing, in 2024 IEEE International Conference on Quantum Computing and Engineering (QCE), Vol.
-1 (IEEE, 2024) pp.
-1315–1323.
 - McPherson [2012] J. W. McPherson, Time dependent dielectric breakdown physics–models revisited, Microelectronics Reliability 52, 1753 (2012).
@@
 D.
-Shelly, Analysis of josephson junction barrier variation – a combined stem, breakdown and monte-carlo approach, arXiv preprint arXiv:2502.13072 (2025).
+Shelly, Analysis of josephson junction barrier variation: A combined electron microscopy, breakdown, and monte carlo approach, Physical Review Materials 9, 084803 (2025).
 - Simmons [1963] J. G. Simmons, Generalized formula for the electric tunnel effect between similar electrodes separated by a thin insulating film, Journal of applied physics 34, 1793 (1963).